Define W =



 x



y







: xy ≥ 0







. Decide if V is a vector space or notand prove



your claim. (Hint: V is the union of the first and third quadrants in the xyplane)

1
Expert's answer
2022-04-19T00:35:36-0400

"\left( 1,2 \right) ,\left( -2,-1 \right) \in V\\\left( 1,2 \right) +\left( -2,-1 \right) =\left( -1,1 \right) \notin V\\Not\,\,a\,\,vector\,\,space"